# ChartAssisstant: A Universal Chart Multimodal Language Model via Chart-to-Table Pre-training and Multitask Instruction Tuning

### 10 Jan 2024 v2

@@ 6 Ablation Study
Ours-D w/o align 87.5 40.4 13.7 35.4 87.9 13.9 52.4 60.3
Ours-D w/o refer 89.2 41.4 14.1 38.6 90.7 14.6 62.2 -
Ours-D 91.3 45.3 14.1 40.2 92.0 14.9 72.1 64.2
The impact of alignment pretraining. We initially validate the importance of alignment pretraining. Table 10 shows that using only multitask instruction tuning falls considerably behind two-stage training strategies. Exact numerical recognition greatly influences mathematical calculation accuracy, leading to a 19.9% and 4.1% performance drop for MathQA and ChartQA-human tasks. Non-mathematical tasks see a 13.9% decrease in OpenCQA compared to the original. We think alignment pre-training, which allows the model to learn chart-table correlations, helps the model better adapt during multitask instruction tuning than handling these processes separately [27].
The impact of referring question answering task. In our experiments, we have observed that integrating referring question answering into multitask instruction tuning training can enhance the model’s performance in other tasks. As shown in table 10, incorporating the referring question answering task leads to improvements across almost all tasks, particularly in tasks requiring mathematical reasoning. For instance, the average performance in ChartQA improves by 3.0%, and in MathQA, it improves by 9.9%. We believe that this task strengthens the model’s ability to understand the visual elements and their relationship in the chart, which contributing to overall performance enhancement [50, 4].
@@ 7 Conclusion
Our work is aimed at developing a generalized multimodal model for chart-related tasks. We propose ChartSFT, a comprehensive and expansive dataset with the most diverse range of supported chart tasks and types. In conjunction, we suggest ChartAssistant, a multimodal model trained using a two-stage strategy over ChartSFT, which can achieve state-of-the-art results across multiple chart-related downstream tasks without task-specific fine-tuning. Through detailed experiments, we further demonstrate the superiority of ChartAssistant.
@@ A ChartSFT
A.1 Numerical QA Templates
We present all the Numerical QA templates in this section. We systematically record both the number of steps in the COT annotation and the number of unique functions used to obtain for each template. Fig 10 shows 101 general templates designed for charts with different types. However, not all of these general templates are applicable to all types of charts. Hence, we’ve customized templates to match the unique characteristics of several specific chart types, such as box plots, bubbles, histograms, and pies, as demonstrated in fig . 15
A.2 Details of Referring QA in ChartSFT
@@ B.1 Experimental Setups
We begin by conducting alignment pre-training, utilizing the chart-to-table translation task for 65k steps. Following that, we engage in multitask instruction tuning. We employ the Adam optimizer [18] with a scheduled learning rate, where the initial rate is set to 5e-5 for ChartAst-D and 2e-6 for ChartAst-S. The input resolution is established at 448×448, while the maximum length in the decoder is defined as 1536 for ChartAst-D and 2048 for ChartAst-S. After training for four epochs for ChartAst-D and only one epoch for ChartAst-S, we perform testing on multiple downstream tasks. During inference, each task receives an image and a textual instruction as input, and the model generates a textual answer. All training processes are carried out on 16xA100 80GB GPUs. ChartAst-S outperforms ChartAst-D and has stronger robustness. This is partly due to the special high-resolution image handling method employed by ChartAst-S, which retains more detailed chart information. Additionally, ChartAst-S incorporates richer pre-training knowledge and the larger model possesses greater robustness.
@@
ChartQA Chart-to-Text Chart-to-Table
Model aug. human Pew Statista ChartQA OpenCQA MathQA ReferQA
stage1 w/o arXiv 89.4 40.8 13.8 39.1 91.0 14.5 64.1 52.3
stage2 w/o arXiv 89.7 40.6 12.7 36.9 91.2 13.2 56.7 47.7
Ours-D 91.3 45.3 14.1 40.2 92.0 14.9 72.1 64.2
@@ C Ablation Study
The impact of arXiv data. We have extracted tabular data from the arXiv dataset, and via multiple APIs, this data has been visualized into suitable charts. Furthermore, it has been expanded into datasets for chart-to-table translation and open-ended question-answering tasks. To validate the importance of this dataset, we conduct experiments by excluding the arXiv data at two distinct stages: the alignment pre-training (stage 1), and the multitask instruction tuning (stage 2).
As shown in table 10, it demonstrates that the arXiv dataset significantly assists the model in aligning charts with tables, thereby improving the performance across various tasks. We believe this is due to the fact that in comparison to existing chart-to-table translation datasets, the arXiv dataset boasts more diversity in terms of style and context; Besides, the open-ended question-answering task contributed by the arXiv dataset is proved to be pivotal for the multitask instruction tuning. We note that the removal of this leads to a drop in the performance of all tasks, most notably math QA and the referring QA. The possible reason for this is because the context and diverse meanings of the arXiv dataset contribute to higher quality question and answering pairs. Therefore, it better promotes multitask tuning.
@@ D Some demos from Out of Distribution
To demonstrate the model’s generalization capability, we randomly take screenshots of several charts, as shown in Fig .7 and Fig .8 . We find that the model possesses generalization ability on out-of-distribution samples.

### 15 Feb 2024 v3

@@ 7 Ablation Study
Ours-D w/o align 89.0 42.1 13.7 38.3 89.5 14.3 62.3 60.1
Ours-D w/o refer 89.2 41.2 14.0 38.6 90.7 14.6 60.2 -
Ours-D 91.3 45.3 14.0 40.2 92.0 14.9 72.1 64.2
The impact of alignment pretraining. We initially validate the importance of alignment pretraining. We ensure that the ”Ours w/o align” version of the model is trained for the same number of iterations as the full ChartAssistant model. Table 7 shows that using only multitask instruction tuning falls considerably behind two-stage training strategies. Exact numerical recognition greatly influences mathematical calculation accuracy, leading to a 9.8% and 3.2% performance drop for MathQA and ChartQA-human tasks. We think alignment pre-training, which allows the model to learn chart-table correlations, helps the model better adapt during multitask instruction tuning than handling these processes separately [27].
The impact of referring question answering task. In our experiments, we have observed that integrating referring question answering into multitask instruction tuning training can enhance the model’s performance in other tasks. As shown in table 7, incorporating the referring question answering task leads to improvements across almost all tasks, particularly in tasks requiring mathematical reasoning. For instance, the average performance in ChartQA improves by 3.1%, and in MathQA, it improves by 11.9%. We believe that this task strengthens the model’s ability to understand the visual elements and their relationship in the chart, which contributing to overall performance enhancement [49, 5].
The impact of arXiv data. we conduct experiments by excluding the arXiv data at two distinct stages: the alignment pre-training (stage 1), and the multitask instruction tuning (stage 2). As shown in table 8, it demonstrates that the arXiv dataset significantly assists the model in aligning charts with tables, thereby improving the performance across various tasks. We believe this is due to the fact that in comparison to existing chart-to-table translation datasets, the arXiv dataset boasts more diversity in terms of style and context; Besides, the open-ended question-answering task contributed by the arXiv dataset is proved to be pivotal for the multitask instruction tuning. We note that the removal of this leads to a drop in the performance of all tasks, most notably math QA and the referring QA. The possible reason for this is because the context and diverse meanings of the arXiv dataset contribute to higher quality question and answering pairs. Therefore, it better promotes multitask tuning.
ChartQA Chart-to-Text Chart-to-Table
Model aug. human Pew Statista ChartQA OpenCQA MathQA ReferQA
stage1 w/o arXiv 89.9 43.7 13.8 39.1 91.1 14.5 64.1 61.1
stage2 w/o arXiv 89.7 42.6 12.6 37.5 91.3 13.2 56.7 56.4
Ours-D 91.3 45.3 14.0 40.2 92.0 14.9 72.1 64.2
@@
Compared with Unichart after task-specific fine-tuning(except for Chart-to-Text). We employ the same training strategy and train with the identical model to highlight the effectiveness gains from our data. Following Unichart’s lead in multitask instruction tuning, as table 9 shows, we fine-tune the model on various test datasets (apart from Chart-to-Text, it utilizes fine-tuning during testing), resulting in improvements across different tasks surpassing those of Unichart. It is noteworthy that both Unichart and ChartAst-D are trained using Donut, emphasizing the superiority of ChartSFT.
Table 9: Compared with Unichart after task-specific fine-tuning.
ChartQA Chart-to-Text Chart-to-Table
Model aug. human Pew Statista ChartQA OpenCQA
Ours-D w/o align 89.0 42.1 13.7 38.3 89.5 14.3
Unichart 87.8 43.9 12.5 38.1 91.1 14.8
Ours-D w/o align(ft) 89.6 44.2 13.7 38.3 91.4 14.9
@@
Key components of ChartSFT analysis. For reasoning tasks involving specific numerical values, such as ChartQA, as shown in table 10, the math question-answering task benefits greatly from this, especially, as illustrated in fig .5, training in COT-format can significantly enhance the accuracy of mathematical computation problems. For tasks involving the output of long texts, such as openCQA, as demonstrated by table 7 and table 8, we find that incorporating a question-answering dataset composed of arXiv data can to some extent improve the performance of these tasks. We believe this is due to the broad scope, diversity, and specificity of the arXiv data. Moreover, compared to SciGraphQA [22], the arXiv data we provide has precise numerical values, results in higher quality question generation. Lastly, thanks to the robust language capabilities of GPT-3.5, it is capable of generating high-quality, comprehensive question-answering datasets.
@@ 8 Conclusion
Our work is aimed at developing a generalized multimodal model for chart-related tasks. We propose ChartSFT, a comprehensive and expansive dataset with the most diverse range of supported chart tasks and types. In conjunction, we suggest ChartAssistant, a multimodal model trained using a two-stage strategy over ChartSFT, which can achieve state-of-the-art results across multiple chart-related downstream tasks. Through detailed experiments, we further demonstrate the superiority of ChartAssistant.
@@
Figure 6: The pipeline of Chart Data Generation in ChartSFT, which consists of three important stages.
@@ A ChartSFT
A.1 Details of Chart Data Generation in ChartSFT
We illustrate the pipeline of data generation in Fig. 6. Concretely, the chart data are generated in the following stages:
Stage 1: Table generation: Taking into account the diversity of tabular data, we have predefined over 20 types of probability density distributions, including normal distribution, uniform distribution, beta distribution, Laplace distribution, and more. For each sample, we randomly choose one type of probability density distribution and utilize it to generate values. For different types of charts, we impose further constraints on these values based on their characteristics. (e.g., value range, ratio of positive and negative values, range interval). For radar, bubble and area charts, We directly utilize randomly generated values as the tabular data. For histogram and box plot, we generate an array of extensive values using this distribution and calculate the statistical metrics of this array to serve as the tabular data (e.g., frequencies corresponding to histograms, upper whiskers corresponding to box plots). And then we use the generated data to prompt ChatGPT for creating titles, legends, and labels that align with the numerical characteristics.
Stage 2: Chart generation: To ensure the diversity of the generated charts, we utilize multiple plot APIs, such as matplotlib, plotly, pyecharts, ggplot, seaborn, altair, and more, to plot a variety of styles of the chart. For each chart, we randomly select the following parameters: line (style, thickness), font (style, size, bold, italic), colors, markers, the position of the elements (title, labels, legends), the size of the charts and so on. Besides our own synthetic tabular data, we also use the table from PlotQA [32], ChartQA [30], ChartSumm [37] and Chart-To-Text [14] to plot the charts for area and radar charts.
Stage 3: Instruction Data generation: For the chart summarization and open-ended QA tasks, we instruct ChatGPT to build datasets by supplying both the table and the corresponding types of charts. For numerical QA and referring QA tasks, we adhere to the approach of the chart with base types by crafting a series of mathematical question templates tailored to the distinct characteristics of various chart types. Subsequently, we manually generate answers with COT annotations.
We adopted a flexible approach by combining ChatGPT with human intervention, which included the utilization of predefined distributions and custom coding of plot API, among other techniques. Through this three-stage chart data generation process, we ensured the diversity and complexity of the table, chart, and instruction data, respectively. As a result, we were able to generate a substantial volume of diversified high-quality chart data.
A.2 Numerical QA Templates
We present all the Numerical QA templates in this section. We systematically record both the number of steps in the COT annotation and the number of unique functions used to obtain for each template. Fig 12 shows 101 general templates designed for charts with different types. However, not all of these general templates are applicable to all types of charts. Hence, we’ve customized templates to match the unique characteristics of several specific chart types, such as box plots, bubbles, histograms, and pies, as demonstrated in fig . 17
A.3 Details of Referring QA in ChartSFT
@@ B.1 Experimental Setups
We begin by conducting alignment pre-training, utilizing the chart-to-table translation task for 65k steps. Following that, we engage in multitask instruction tuning. We employ the Adam optimizer [17] with a scheduled learning rate, where the initial rate is set to 5e-5 for ChartAst-D and 2e-6 for ChartAst-S. The input resolution is established at 224×224 and 448×448, while the maximum length in the decoder is defined as 1536 for ChartAst-D and 2048 for ChartAst-S. After training for four epochs for ChartAst-D and only one epoch for ChartAst-S, we perform testing on multiple downstream tasks. During inference, each task receives an image and a textual instruction as input, and the model generates a textual answer. All training processes are carried out on 16xA100 80GB GPUs. ChartAst-S outperforms ChartAst-D and has stronger robustness. This is partly due to the special high-resolution image handling method employed by ChartAst-S, which retains more detailed chart information. Additionally, ChartAst-S incorporates richer pre-training knowledge and the larger model possesses greater robustness.
@@ D Some demos from Out of Distribution
To demonstrate the model’s generalization capability, we randomly take screenshots of several charts, as shown in Fig .9 and Fig .10 . We find that the model possesses generalization ability on out-of-distribution samples. Additionally, as shown in fig. 8, we visualize some demos comparing the performance of zero-shot scenarios with baseline methods. We observe that in summarization tasks, UniChart and Matcha tend to produce repetitions or hallucinations, whereas ChartLlama and ChartAssistant exhibit relatively stronger capabilities in handling summarization tasks. However, ChartLlama commits some factual errors; in question answering, thanks to the incorporation of COT-format QA training data, ChartAssistant effectively addresses QA tasks requiring mathematical reasoning. Lastly, in chart-to-table translation, UniChart and Matcha accurately model the table structure. Although ChartLlama can model the table structure accurately, the values are completely incorrect. Only ChartAssistant successfully constructs the table of the chart accurately.